# UAV-Rain1k: A Benchmark for Raindrop Removal from UAV Aerial Imagery

### v2

@@ 3.1 Quantitative Evaluation
It is evident from the results that Restormer [23] and DRSformer [3] achieve the top two quantitative outcomes in rain removal performance. However, it is noteworthy that their model complexity is comparatively higher than traditional CNN methods [16, 22]. To comprehensively evaluate the performance and efficiency of different algorithms, future research endeavors could delve deeper into strategies for maintaining high performance while reducing model complexity. This would address the requirements of speed and resource consumption in real-world remote sensing applications [2].
@@ 3.2 Qualitative Evaluation
Figure 6 and Figure 7 illustrates the visual comparison results of different baseline algorithms on our proposed benchmark (In Figure 6, the effect corresponds to a composite image taken from a horizontal angle, while Figure 7 represents a composite image captured from an downward angle). In comparison to CNN-based methods (RCDNet [16] and SPDNet [22]), Transformer-based approaches exhibit superior capability in removing undesired raindrop and recovering clearer visual outcomes, it achieves remarkable visual effects from any aerial perspective, eliminating raindrops and obscuring effects without rain residues and pseudo artifacts. owing to the evident advantage of Transformers in modeling non-local information. Additionally, we observe grid artifacts in the restoration results of IDT [19] and DRSformer [3], indicating their limited capacity in handling high-resolution UAV imagery. Furthermore, nearly all methods exhibit poor performance in local detail regions, as they tend to remove some image details while removing raindrop, suggesting that there are still rooms for improvement.
@@ 4 Conclusion
In this paper, we have proposed a new benchmark dataset for raindrop removal from UAV aerial imagery. We have provided a detailed overview of the synthesis process for the UAVRain1k dataset, covering background collection, raindrop modeling, image synthesis, results demonstration, and subjective evaluation. Based on the proposed new benchmark, we further experimentally validated representative image deraining algorithms in three aspects: quantitative evaluation, qualitative evaluation, and application-based evaluation. The experimental results shed light on the limitations of existing methods and point towards promising future directions. This paper not only pioneers a new research direction in UAV image processing but also advocates for the community to propose more effective algorithms.

### v3

@@ 3.1 Quantitative Evaluation
It is evident from the results that DRSformer [3] and Restormer [23] achieve the top two quantitative outcomes in rain removal performance. However, it is noteworthy that their model complexity is comparatively higher than traditional CNN methods [16, 22]. To comprehensively evaluate the performance and efficiency of different algorithms, future research endeavors could delve deeper into strategies for maintaining high performance while reducing model complexity. This would address the requirements of speed and resource consumption in real-world remote sensing applications [2].
@@ 3.2 Qualitative Evaluation
Figure 6 and Figure 7 illustrates the visual comparison results of different baseline algorithms on our proposed benchmark (In Figure 6, the effect corresponds to a composite image taken from a horizontal angle, while Figure 7 represents a composite image captured from an downward angle). In comparison to CNN-based methods (RCDNet [16] and SPDNet [22]), Transformer-based approaches exhibit superior capability in removing undesired raindrop and recovering clearer visual outcomes, it achieves remarkable visual effects from any aerial perspective, eliminating raindrops and obscuring effects without rain residues and pseudo artifacts. owing to the evident advantage of Transformers in modeling non-local information. Additionally, we observe grid artifacts in the restoration results of IDT [19], indicating their limited capacity in handling high-resolution UAV imagery. Furthermore, nearly all methods exhibit poor performance in local detail regions, as they tend to remove image details while removing raindrop, suggesting that there are still rooms for improvement.
@@ 4 Conclusion
In this paper, we have proposed a new benchmark dataset for raindrop removal from UAV aerial imagery. We have provided a detailed overview of the synthesis process for the UAV-Rain1k dataset, covering background collection, raindrop modeling, image synthesis, results demonstration, and subjective evaluation. Based on the proposed benchmark, we further experimentally validated representative image deraining algorithms in three aspects: quantitative evaluation, qualitative evaluation, and application-based evaluation. Experimental results shed light on the limitations of existing methods and point towards promising future directions. This paper not only pioneers a new research direction in UAV image processing but also advocates for the community to propose more effective algorithms.